Question 1

Maman-Paz is a distributor of meal kits. Their meal kit is a package that contains all necessary food items for customers cook it at home. They source ingredients from manufacturers and distribute to retailers such as Walmart and Target. To date they have not been reaching the Boston area and had no customers from there. However, with new budget approved they are ready to enter this market.

Maman-Paz has selected 6 target cities they want to serve with their famous pasta kit, PastaFusion, in the first year of their operations. To serve these cities they need to purchase warehouses from a number of candidate locations and plan inbound deliveries from suppliers to warehouses as well as outbound deliveries from warehouses to the target cities.

The table below shows the target cities and their forecasted demand for PastaFusion the coming year.

Cambridge: 10 Tons/year
Medford: 7 Tons/year
Boston: 12 Tons/year
Quincy: 3 Tons/year
Winthrope: 9 Tons/year
Dorchester: 4 Tons/year

Four candidate warehouse locations have been identified in the following areas with their estimated capacity.

Allston: 35 Tons/year
Somerville: 60 Tons/year
Chelsea: 20 Tons/year
Malden: 40 Tons/year

Maman-Paz potential suppliers are located in the following locations with the following capacities:

Arlington: 35 Tons/year
Brookline: 40 Tons/year
Melrose: 25 Tons/year

Below are the distance matrices, in miles, from suppliers to warehouses and from warehouses to target cities. Shipping cost from suppliers to warehouses is $0.98 per Ton per mile. Shipping cost from warehouses to target cities is $1.57 per Ton per mile.

Allston

Somerville

Chelsea

Malden

Arlington

130

110

100

150

Brookline

70

80

90

60

Melrose

120

130

130

100

Cambridge

Medford

Boston

Quincy

Winthrope

Dorchester

Allston

10

30

30

50

100

50

Somerville

20

40

15

40

110

40

Chelsea

50

40

5

50

100

20

Malden

10

5

50

90

65

40

They ask you to solve the network design problem. Management also notifies you that it is not necessary to use all of the potential suppliers.

Once you present your solutions and receive feedback, you discover that maintaining warehouses also has a cost and this cost must be considered in the model. This cost is proportional to the warehouse capacity. For every Ton per year of capacity, there is $10000 maintenance cost. For example, if a warehouse’s capacity is 40 Tons/year, the company incurs $400000 maintenance cost per year.

Assuming the original warehouse capacities, add the new maintenance cost to the model and solve the problem again.

For those warehouses that are open, what percentage of the total available warehouse capacity is being used?

Enter a value between 0 and 1 with TWO decimal places.
Answer #1

1)

Optimal solution is determined using Solver as follows:

xSUMPRODUCT(B4:E6,J4:M6)*0.98+SUMPRODUCT(B10:G13,J10:013) 1.57 U18 Tables of optimal distribution plan Parameters Tables Alls

EXCEL FORMULAS:

Tables of optimal distribution plan
Allston Somerville Chelsea Malden Row Total
Arlington 0 0 5 0 =SUM(J4:M4)
Brookline 0 3 11 26 =SUM(J5:M5)
Melrose 0 0 0 0 =SUM(J6:M6)
Column Total =SUM(J4:J6) =SUM(K4:K6) =SUM(L4:L6) =SUM(M4:M6)
Cambridge Medford Boston Quincy Winthrope Dorchester Row Total
Allston 0 0 0 0 0 0 =SUM(J10:O10)
Somerville 0 0 0 3 0 0 =SUM(J11:O11)
Chelsea 0 0 12 0 0 4 =SUM(J12:O12)
Malden 10 7 0 0 9 0 =SUM(J13:O13)
Column Total =SUM(J10:J13) =SUM(K10:K13) =SUM(L10:L13) =SUM(M10:M13) =SUM(N10:N13) =SUM(O10:O13)

Total cost =SUMPRODUCT(B4:E6,J4:M6)*0.98+SUMPRODUCT(B10:G13,J10:O13)*1.57

Solver parameters:

SSolver Parameters SJS18 Set Objective O Max Min O Value Of: o Of: 0 By Changing aiable Cells: SJS4:SMS6,SJS10:SOS13 Subject

2)

The modified model and solution is following:

e fx SUMPRODUCT(B4:E6,J4:M6) *0.98+SUMPRODUCT(B11:G14,J11:014) 1.57+SUM(B8:E8) 10000 J19 Parameters Tables Tables of optimal

EXCEL FORMULAS:

Parameters Tables
Allston Somerville Chelsea Malden Capacity
Arlington 130 110 100 150 35
Brookline 70 80 90 60 40
Melrose 120 130 130 100 25
Capacity 35 60 20 40
Logical Capacity =B7*J8 =C7*K8 =D7*L8 =E7*M8
Cambridge Medford Boston Quincy Winthrope Dorchester
Allston 10 30 30 50 100 50
Somerville 20 40 15 40 110 40
Chelsea 50 40 5 50 100 20
Malden 10 5 50 90 65 40
Demand 10 7 12 3 9 4

Total cost (Cell J19)  =SUMPRODUCT(B4:E6,J4:M6)*0.98+SUMPRODUCT(B11:G14,J11:O14)*1.57+SUM(B8:E8)*10000

Solver parameters:

SSolver Parameters SIS19 Set Objective O Max ⓔMin O Value Of: 0 By Changing aiable Cells: SJS4:SMS6,SJS11:SOS14,SJS8:SMS8 Sub

Allston and Chelsea warehouses should be kept open

Percentage of total available warehouse capacity used of Allston = 29/35 = 83 %

Percentage of total available warehouse capacity used of Chelsea = 16/20 = 80 %

Total cost = $ 555,805